# Optimizing Treatment Allocation in the Presence of Interference

### v1

@@ Abstract
In Influence Maximization (IM), the objective is to – given a budget – select the optimal set of entities in a network to target with a treatment so as to maximize the total effect. For instance, in marketing, the objective is to target the set of customers that maximizes the total response rate, resulting from both direct treatment effects on targeted customers and indirect, spillover, effects that follow from targeting these customers. Recently, new methods to estimate treatment effects in the presence of network interference have been proposed. However, the issue of how to leverage these models to make better treatment allocation decisions has been largely overlooked. Traditionally, in Uplift Modeling (UM), entities are ranked according to estimated treatment effect, and the top entities are allocated treatment. Since, in a network context, entities influence each other, the UM ranking approach will be suboptimal. The problem of finding the optimal treatment allocation in a network setting is combinatorial and generally has to be solved heuristically. To fill the gap between IM and UM, we propose OTAPI: Optimizing Treatment Allocation in the Presence of Interference to find solutions to the IM problem using treatment effect estimates. OTAPI consists of two steps. First, a causal estimator is trained to predict treatment effects in a network setting. Second, this estimator is leveraged to identify an optimal treatment allocation by integrating it into classic IM algorithms. We demonstrate that this novel method outperforms classic IM and UM approaches on both synthetic and semi-synthetic datasets.
@@ 1 Introduction
A well-studied and important problem within network analysis is Influence Maximization (IM) (Kempe et al., 2003; Domingos & Richardson, 2001; Richardson & Domingos, 2002; Azaouzi et al., 2021). The primary goal in IM is to identify the set of entities in a network to target with a particular intervention, e.g., a vaccine or a marketing campaign, in order to maximize the intervention’s effect across the entire network given a budget $k$. The effect of treating one entity can cascade through the network, affecting connected entities as well. Consequently, careful analysis is essential to allocate treatment efficiently. For instance, in viral marketing, (potential) customers influence each other’s purchasing decisions (Goldenberg et al., 2001). Hence, when determining whom to target with a marketing campaign, merely estimating the Individual Treatment Effects (ITEs), as commonly done in Uplift Modeling (UM) (Devriendt et al., 2018), may not suffice. Considering interference effects – where treating one entity may influence others – is critical to identify the optimal set of entities to target. Practical applications of this problem include, e.g., vaccination and marketing campaigns.
@@
Influence Maximization.  The IM problem can be formalized as follows: given a network $\mathcal{G}=(\mathcal{N},\mathcal{E})$, where $\mathcal{N}$ represents the set of nodes and $\mathcal{E}$ denotes the set of edges, the objective is to identify the set of $k$ seed nodes that are to be treated in order to maximize the total influence spread, i.e., the total effect of treatment allocation (Kempe et al., 2003). In the IM literature, the influence spread of a certain treatment allocation is typically modeled using a diffusion process. This process describes how information or behavior disseminates through the network (Granovetter, 1978; Goldenberg et al., 2001).
@@
We identify three significant limitations in the current approaches to address the IM problem. Firstly, node features are mostly ignored. Nevertheless, these features may contain important information regarding the spread of behaviors or effects through the network (Ogburn & VanderWeele, 2014; Ma & Tresp, 2021). Secondly, to the best of our knowledge, existing methods typically assume a homogeneous effect of treatment. More specifically, it is assumed that all entities will respond positively (e.g., purchase the product) if they receive treatment. Yet, in many applications, treatment effects vary among entities, and the susceptibility to treatment often depends on entity characteristics (Shalit et al., 2017; Athey & Imbens, 2016). Lastly, the most commonly used diffusion processes rely on assumptions rather than data. However, employing an ill-defined diffusion process to solve the IM problem may lead to suboptimal outcomes (Aral & Dhillon, 2018).
@@
Uplift Modeling.  Another method for optimizing treatment allocation is UM (Devriendt et al., 2018). This approach involves estimating the ITE for each entity, ranking them according to these estimates, and selecting the top $k$ ranked entities with the highest expected ITE for treatment. An important limitation of this approach, however, is its assumption that entities are independent of each other. This assumption is violated in network settings where entities influence each other, such as marketing or vaccination campaigns. Recently, advanced causal inference methods have been proposed to cope with such network effects, known as spillover effects, by incorporating network information to estimate treatment effects (Forastiere et al., 2021; Jiang & Sun, 2022; Cai et al., 2023; Huang et al., 2023). However, even with the capability to estimate treatment effects in the presence of interference, determining the optimal treatment allocation has been largely overlooked. Ranking entities according to their estimated treatment effects is no longer a possibility because the decision to treat one entity may impact the decision whether or not to treat another entity.
@@
Figure 1: Overview of OTAPI. In Step 1, a relational causal estimator is trained on historical data. It receives the treatment assignment $\mathbf{t}$ and the network $\mathcal{G}$, including the features $\mathbf{x}_{i}$ for each node $i$ (represented by the colors) as input. The estimator outputs the predicted outcomes $\hat{y}_{i}$ for each node. To train the model, these predictions are compared with the actual outcomes $y_{i}$. In Step 2, an optimization algorithm (e.g., a genetic algorithm) is used to find the optimal treatment allocation $\mathbf{t^{*}}$ for a potentially unseen network (e.g., a network that has changed over time). This optimization algorithm takes as input the new network and the set of feasible allocations, which depends on the budget $k$. The trained causal model is leveraged to estimate the total effect of treatment allocation, which is used as the objective function in the optimization algorithm.
@@
OTAPI.  We propose OTAPI: Optimizing Treatment Allocation in the Presence of Interference, a novel method that combines the strengths of both UM and IM approaches to identify improved, data-driven, solutions to the IM problem. Our method consists of two steps. First, a relational causal estimator is trained using observational data, e.g., from historical marketing or vaccination campaigns. A relational causal estimator is an estimator that can leverage both individual and network information to estimate causal effects. This estimator allows estimating the total effect of any treatment allocation and controls for confounding. Second, the treatment effect estimates from this causal estimator are utilized to optimize the treatment allocation. Given the combinatorial nature of this optimization problem, this step presents a significant challenge. To address this challenge, treatment allocation is optimized by integrating treatment effect estimates into existing IM algorithms. An overview of OTAPI is depicted in Figure 1.
@@
Contributions.  To the best of our knowledge, the challenge of optimally allocating treatment in the presence of interference is yet to be addressed. We conjecture that leveraging relational causal estimators will aid in finding the optimal treatment allocation. To this end, (1) we formalize the problem of finding the optimal treatment allocation in the presence of interference; (2) we present OTAPI, a new two-step method for optimizing treatment allocation in the presence of interference; and (3) we run extensive experiments on both synthetic and semi-synthetic data to empirically demonstrate that OTAPI finds better treatment allocations (in terms of the total effect of treatment allocation) than traditional methods used in UM and IM.
@@ 2.1 Treatment effect estimation in the presence of interference
A substantial body of literature has focused on treatment effect estimation when entities are independent (Johansson et al., 2016; Yao et al., 2018; Devriendt et al., 2018). Within this literature, an important and extensively studied problem is confounding bias (Rosenbaum & Rubin, 1983; Shalit et al., 2017; Vairetti et al., 2024). This bias is associated with the treatment assignment mechanism, whereby treatment is not assigned randomly. Consequently, the distributions of the control and treatment groups will differ. In most applications, some form of mechanism or policy typically governs treatment assignment, leading to potential bias in treatment effect estimates. For instance, in marketing, a company might target a specific subgroup of the population. As it may be costly or even impossible to randomize the treatment assignment, (potentially biased) observational data has to be used.
Recent approaches construct balanced representations of the distributions of the control and treatment groups to mitigate confounding bias (Shalit et al., 2017; Johansson et al., 2016; Yao et al., 2018; Hassanpour & Greiner, 2019). In a network setting, where entities can influence each other, mitigating confounding bias becomes more challenging. The features and treatments of connected entities might influence the probability of receiving treatment, introducing confounding bias as well (Ma & Tresp, 2021; Jiang & Sun, 2022).
In estimating treatment effects in the presence of interference, a common approach involves defining an exposure mapping (Aronow & Samii, 2017), which generates a variable (or set of variables) based on the treatments of other relevant entities. Many methods assume this exposure mapping to be the percentage of directly connected entities, known as neighbors, that received treatment and use Graph Neural Networks (GNNs) (Kipf & Welling, 2016; Xu et al., 2018) to learn a model from observational data (Ma & Tresp, 2021; Jiang & Sun, 2022; Cai et al., 2023). Representation balancing techniques are employed to mitigate confounding bias. Some recent works propose to relax the assumption on the exposure mapping to allow for heterogeneous peer influence, where the spillover effect may vary depending on the features of connected entities (Huang et al., 2023; Adhikari & Zheleva, 2023). Another line of work focuses on leveraging network information, such as features of neighbors, to better control for hidden confounders. For example, a person’s community affiliations may help to represent their socioeconomic status. However, these methods typically ignore spillover effects, assuming that the treatment of one entity does not causally affect other entities (Guo et al., 2020, 2021).
@@ 2.2 Influence Maximization problem
The IM problem was initially introduced by Domingos & Richardson (2001), and later formalized as a combinatorial optimization problem by Kempe et al. (2003). A multitude of methods to tackle the IM problem rely on diffusion processes, which define how influence propagates through the network. A wide variety of diffusion processes have been proposed, each with its distinct assumptions about the dissemination of behavior or information (Goldenberg et al., 2001; Granovetter, 1978; Kempe et al., 2003; Ye et al., 2022). To identify the optimal allocation set, existing methods mostly employ a greedy algorithm proposed by Kempe et al. (2003). This algorithm uses Monte Carlo (MC) simulation based on the diffusion process to estimate the expected influence spread, i.e., the expected number of activated entities. However, its computational complexity may render it infeasible in terms of runtime, especially when networks become large. To this end, several methods have been developed to reduce the number of MC simulations, such as CELF (Leskovec et al., 2007) and CELF++ (Goyal et al., 2011). Other methods were developed to partition the network into communities, significantly reducing the runtime of an MC simulation (Wang et al., 2010; Bozorgi et al., 2017). Alternatively, some approaches ignore diffusion processes altogether, instead solely using the network structure. These methods employ, for instance, the degree centrality, which is defined as the number of connections of a node. IM algorithms based on this measure include the degree (Kempe et al., 2003) and single discount heuristics (Chen et al., 2009).
@@
All the aforementioned methods rely on the network structure or make assumptions regarding the diffusion process in order to arrive at a solution. It has been shown, however, that incorporating data to estimate the spread of influence through the network yields more reliable solutions (Aral & Dhillon, 2018). In the IM literature, considerable work has been done to employ propagation traces to estimate the parameters of specific diffusion processes (Goyal et al., 2010; Saito et al., 2008; Bourigault et al., 2016; Tran & Zheleva, 2022). Propagation traces are data structures containing information about which entities were activated and when. Goyal et al. (2011) uses these traces to create a model that does not make any parametric assumptions about the diffusion process, but instead learns the expected influence spread directly from data. These methods, however, do not consider node features. The only exception is Tran & Zheleva (2022), in which node features are leveraged to estimate the parameters in the Linear Threshold model, a commonly used diffusion process. In contrast to these approaches, our proposed method leverages node features, does not make parametric assumptions regarding the diffusion process, and does not use propagation traces. Moreover, OTAPI allows for out-of-sample IM, as the network for which we try to find the optimal treatment allocation may be different from the network that was used for training the model. This is an important advantage of OTAPI, as networks – especially social networks – typically change over time. Ma & Tresp (2021) proposes a utility maximization approach for instance-level treatment assignment in the presence of interference using a policy network that leverages a relational causal estimator. The budget constraint is satisfied by integrating it into the loss function of the policy network and allowing some tolerance for violating it. In comparison, OTAPI directly optimizes treatment allocation for the network as a whole by leveraging algorithms that can support hard constraints.
@@ 3 Problem setup
Notation.  Consider an undirected network $\mathcal{G}=(\mathcal{N},\mathcal{E})$ where $\mathcal{N}$ represents the set of nodes and $\mathcal{E}$ denotes the set of edges between them. We define the set of neighbors of node $i$ as $\mathcal{N}_{i}$, which comprises nodes $j\in\mathcal{N}$ for which an edge $(i,j)$ exists. In this context, $(i,j)=(j,i)$. Additionally, we denote the set of nodes not in the neighborhood of $i$ as $\mathcal{N}_{-i}=\mathcal{N}\setminus(\mathcal{N}_{i}\cup\{i\})$. Each node $i$ possesses a feature vector $\mathbf{X}_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$. We will use nodes and entities interchangeably, depending on the context. An entity is defined by an index $i$ and has a feature vector $\mathbf{X}_{i}$.
Let $Y_{i}\in\mathcal{Y}$ represent the outcome of entity $i$. Without loss of generality, $\mathcal{Y}=\{0,1\}$, e.g., buying or not buying the product. We define $T_{i}\in\mathcal{T}=\{0,1\}$ as the treatment given to entity $i$. The treatments of $i$’s neighbors $\mathcal{N}_{i}$ is given by $\mathbf{T}_{\mathcal{N}_{i}}=\{T_{j}\}_{j\in\mathcal{N}_{i}}$. Following (Forastiere et al., 2021; Ma & Tresp, 2021; Aronow & Samii, 2017), an exposure mapping, denoted by $Z:\mathbf{T}_{\mathcal{N}_{i}}\to Z_{i}$, is defined to summarize $\mathbf{T}_{\mathcal{N}_{i}}$. This function describes how the outcome of an entity is dependent on $\mathbf{T}_{\mathcal{N}_{i}}$. The domain $\mathcal{Z}$ of the exposure mapping $Z_{i}$ depends on the exact definition of the function $Z$ (Forastiere et al., 2021).
@@
Figure 2: Graphical representation of the causal structure. Only two entities are included for simplicity. $Y_{i}$ is influenced by its own features $\mathbf{X}_{i}$, the features of its neighbors $\mathbf{X}_{j}$, its own treatment $T_{i}$, and the treatment of its neighbors $T_{j}$. Moreover, $T_{i}$ is influenced by both $\mathbf{X}_{i}$ and $\mathbf{X}_{j}$
Assumptions.  We assume that an entity’s own treatment $T_{i}$, the treatments of its neighbors $\mathbf{T}_{\mathcal{N}_{i}}$, an entity’s features $\mathbf{X_{i}}$ and the features of its neighbors $\mathbf{X}_{\mathcal{N}_{i}}=\{\mathbf{X}_{j}\}_{j\in\mathcal{N}_{i}}$ causally influence its outcome $Y_{i}$. This causal structure is depicted in Figure 2. Note that we do not consider contagion effects, meaning the outcomes of different entities do not causally affect each other (Ogburn & VanderWeele, 2014). To summarize $\mathbf{T}_{\mathcal{N}_{i}}$ we use the mapping $Z_{i}=\frac{\sum_{j\in\mathcal{N}_{i}}T_{j}}{|\mathcal{N}_{i}|}$ (Forastiere et al., 2021; Jiang & Sun, 2022). Note that more complex exposure mappings (e.g., multi-dimensional, based on node features, etc.), can also be defined/learned (Cai et al., 2023; Adhikari & Zheleva, 2023; Aronow & Samii, 2017). We assume that the decision to treat is taken simultaneously for all entities. Consequently, both $Z_{i}$ and $T_{i}$ are part of the intervention on $i$.
Traditionally, each entity has two potential outcomes: $Y_{i}(T_{i}=0)$ and $Y_{i}(T_{i}=1)$. However, in the presence of interference, the traditional consistency assumption, which states that there are no multiple versions of treatment, is violated. Here, depending on the treatment of $i$’s neighbors, $i$’s potential outcomes might differ. To address this, the consistency assumption is relaxed by employing the exposure mapping $Z_{i}$ (Forastiere et al., 2021; Jiang & Sun, 2022):

### v2

@@ Abstract
In Influence Maximization (IM), the objective is to—given a budget—select the optimal set of entities in a network to target with a treatment so as to maximize the total effect. For instance, in marketing, the objective is to target the set of customers that maximizes the total response rate, resulting from both direct treatment effects on targeted customers and indirect, spillover, effects that follow from targeting these customers. Recently, new methods to estimate treatment effects in the presence of network interference have been proposed. However, the issue of how to leverage these models to make better treatment allocation decisions has been largely overlooked. Traditionally, in Uplift Modeling (UM), entities are ranked according to estimated treatment effect, and the top entities are allocated treatment. Since, in a network context, entities influence each other, the UM ranking approach will be suboptimal. The problem of finding the optimal treatment allocation in a network setting is NP-hard, and generally has to be solved heuristically. To fill the gap between IM and UM, we propose OTAPI: Optimizing Treatment Allocation in the Presence of Interference to find solutions to the IM problem using treatment effect estimates. OTAPI consists of two steps. First, a causal estimator is trained to predict treatment effects in a network setting. Second, this estimator is leveraged to identify an optimal treatment allocation by integrating it into classic IM algorithms. We demonstrate that this novel method outperforms classic IM and UM approaches on both synthetic and semi-synthetic datasets.
@@ 1 Introduction
A well-studied and important problem within network analysis is Influence Maximization (IM) (Kempe et al., 2003; Domingos and Richardson, 2001; Richardson and Domingos, 2002; Azaouzi et al., 2021). The primary goal in IM is to identify the set of entities in a network to target with a particular intervention, e.g., a vaccine or a marketing campaign, in order to maximize the intervention’s effect across the entire network given a budget $k$. The effect of treating one entity can cascade through the network, affecting connected entities as well. Consequently, careful analysis is essential to allocate treatment efficiently. For instance, in viral marketing, (potential) customers influence each other’s purchasing decisions (Goldenberg et al., 2001). Hence, when determining whom to target with a marketing campaign, merely estimating the Individual Treatment Effects (ITEs), as commonly done in Uplift Modeling (UM) (Devriendt et al., 2018), may not suffice. Considering interference effects—where treating one entity may influence others—is critical to identify the optimal set of entities to target. Practical applications of this problem include, e.g., vaccination and marketing campaigns.
@@
Influence Maximization.  The IM problem can be formalized as follows: given a network $\mathcal{G}=(\mathcal{N},\mathcal{E})$, where $\mathcal{N}$ represents the set of nodes and $\mathcal{E}$ denotes the set of edges, the objective is to identify the set of $k$ seed nodes that are to be treated in order to maximize the total influence spread, i.e., the total effect of treatment allocation. This problem has been shown to be NP-hard (Kempe et al., 2003). In the IM literature, the influence spread of a certain treatment allocation is typically modeled using a diffusion process. This process describes how information or behavior disseminates through the network (Granovetter, 1978; Goldenberg et al., 2001).
@@
We identify three significant limitations in the current approaches to address the IM problem. First, node features are mostly ignored. Nevertheless, these features may contain important information regarding the spread of behaviors or effects through the network (Ogburn and VanderWeele, 2014; Ma and Tresp, 2021). Second, existing methods typically assume a homogeneous direct effect of treatment. More specifically, it is assumed that all entities will respond positively (e.g., purchase the product) if they receive treatment. Yet, in many applications, treatment effects vary among entities, and the susceptibility to treatment often depends on entity characteristics (Shalit et al., 2017; Athey and Imbens, 2016). Last, the most commonly used diffusion processes rely on assumptions rather than data. However, employing an ill-defined diffusion process to solve the IM problem may lead to suboptimal outcomes (Aral and Dhillon, 2018).
@@
Uplift Modeling.  Another method for optimizing treatment allocation is UM (Devriendt et al., 2018). This approach involves estimating the ITE for each entity, ranking them according to these estimates, and selecting the top $k$ ranked entities with the highest expected ITE for treatment. An important limitation of this approach, however, is its assumption that entities are independent of each other. This assumption is violated in network settings where entities influence each other, such as marketing or vaccination campaigns. Recently, advanced causal inference methods have been proposed to cope with such network effects, known as spillover effects, by incorporating network information to estimate treatment effects (Forastiere et al., 2021; Jiang and Sun, 2022; Cai et al., 2023; Huang et al., 2023). However, even with the capability to estimate treatment effects in the presence of interference, determining the optimal treatment allocation has been largely overlooked. Ranking entities according to their estimated treatment effects is no longer a possibility because the decision to treat one entity may impact the decision whether or not to treat another entity.
@@
Figure 1: Overview of OTAPI. In Step 1, a relational causal estimator is trained on historical data. It receives the treatment assignment $\mathbf{t}$ and the network $\mathcal{G}$, including the features $\mathbf{x}_{i}$ for each node $i$ (represented by the colors) as input. The estimator outputs the predicted outcomes $\hat{y}_{i}$ for each node. To train the model, these predictions are compared with the actual outcomes $y_{i}$. In Step 2, an optimization algorithm (e.g., a genetic or greedy algorithm) is used to find the optimal treatment allocation $\mathbf{t^{*}}$ for a potentially unseen network (e.g., a network that has changed over time). This optimization algorithm takes as input the new network and the set of feasible allocations, which depends on the budget $k$. The trained causal model is leveraged to estimate the total effect of treatment allocation, which is used as the objective function in the optimization algorithm.
@@
OTAPI.  We propose OTAPI: Optimizing Treatment Allocation in the Presence of Interference, a novel method that combines the strengths of both UM and IM approaches to identify improved, data-driven, solutions to the IM problem. Our method consists of two steps. First, a relational causal estimator is trained using observational data, e.g., from historical marketing or vaccination campaigns. A relational causal estimator is an estimator that can leverage both individual and network information to estimate causal effects. This estimator allows estimating the total effect of any treatment allocation and controls for confounding. Second, the treatment effect estimates from this causal estimator are integrated into existing IM algorithms in order to optimize the treatment allocation. An overview of OTAPI is depicted in Figure 1.
@@
Contributions.The challenge of data-driven treatment allocation that accounts for interference is yet to be addressed. We conjecture that leveraging relational causal estimators will aid in finding the optimal treatment allocation. To this end, (1) we formalize the problem of finding the optimal treatment allocation in the presence of interference; (2) we present OTAPI, a new two-step method for optimizing treatment allocation in the presence of interference; and (3) we run extensive experiments on both synthetic and semi-synthetic data to empirically demonstrate that OTAPI finds better treatment allocations (in terms of the total effect of treatment allocation) than traditional methods used in UM and IM.
@@ 2.1 Treatment effect estimation in the presence of interference
A substantial body of literature has focused on treatment effect estimation when entities are independent (Johansson et al., 2016; Yao et al., 2018; Devriendt et al., 2018). Within this literature, an important and extensively studied problem is confounding bias, also sometimes referred to as the covariate shift problem (Rosenbaum and Rubin, 1983; Shalit et al., 2017; Vairetti et al., 2024). This bias is associated with the treatment assignment mechanism, whereby treatment is not assigned randomly. Consequently, the distributions of the control and treatment groups will differ. In most applications, some form of mechanism or policy typically governs treatment assignment, leading to potential bias in treatment effect estimates. For instance, in marketing, a company might target a specific subgroup of the population. As it may be costly or even impossible to randomize the treatment assignment, (potentially biased) observational data has to be used.
Recent approaches construct balanced representations of the distributions of the control and treatment groups to mitigate confounding bias (Shalit et al., 2017; Johansson et al., 2016; Yao et al., 2018; Hassanpour and Greiner, 2019). In a network setting, where entities can influence each other, mitigating confounding bias becomes more challenging. The features and treatments of connected entities might influence the probability of receiving treatment, introducing confounding bias as well (Ma and Tresp, 2021; Jiang and Sun, 2022).
In estimating treatment effects in the presence of interference, a common approach involves defining an exposure mapping (Aronow and Samii, 2017), which generates a variable (or set of variables) based on the treatments of other relevant entities. Many methods assume this exposure mapping to be the percentage of directly connected entities, known as neighbors, that received treatment, and use Graph Neural Networks (GNNs) (Kipf and Welling, 2016; Xu et al., 2018) to learn a model from observational data (Ma and Tresp, 2021; Jiang and Sun, 2022; Cai et al., 2023). Representation balancing techniques are employed to mitigate confounding bias. Some recent works propose to relax the assumption on the exposure mapping to allow for heterogeneous peer influence, where the spillover effect may vary depending on the features of connected entities (Huang et al., 2023; Adhikari and Zheleva, 2023; Zhao et al., 2024). Another line of work focuses on leveraging network information, such as features of neighbors, to better control for hidden confounders. For example, a person’s community affiliations may help to represent their socioeconomic status. However, these methods typically ignore spillover effects, assuming that the treatment of one entity does not causally affect other entities (Guo et al., 2020, 2021).
@@ 2.2 Influence Maximization problem
The IM problem was initially introduced by Domingos and Richardson (2001), and later formalized as a combinatorial optimization problem by Kempe et al. (2003). A multitude of methods to tackle the IM problem rely on diffusion processes, which define how influence propagates through the network. A wide variety of diffusion processes have been proposed, each with its distinct assumptions about the dissemination of behavior or information (Goldenberg et al., 2001; Granovetter, 1978; Kempe et al., 2003; Ye et al., 2022). To identify the optimal allocation set, existing methods mostly employ a greedy algorithm proposed by Kempe et al. (2003). This algorithm uses Monte Carlo (MC) simulations based on the diffusion process to estimate the expected influence spread, i.e., the expected number of activated entities. However, its computational complexity may render it infeasible in terms of runtime, especially when networks become large. To this end, several methods have been developed to reduce the number of MC simulations, such as CELF (Leskovec et al., 2007) and CELF++ (Goyal et al., 2011b). Other methods were developed to partition the network into communities, significantly reducing the runtime of an MC simulation (Wang et al., 2010; Bozorgi et al., 2017). Alternatively, some approaches ignore diffusion processes altogether, instead solely using the network structure. These methods employ, for instance, the degree centrality, which is defined as the number of connections of a node. IM algorithms based on this measure include the degree (Kempe et al., 2003) and single discount heuristics (Chen et al., 2009).
@@
All the aforementioned methods rely on the network structure or make assumptions regarding the diffusion process in order to arrive at a solution. It has been shown, however, that incorporating data to estimate the spread of influence through the network yields more reliable solutions (Aral and Dhillon, 2018). In the IM literature, considerable work has been done to employ propagation traces to estimate the parameters of specific diffusion processes (Goyal et al., 2010; Saito et al., 2008; Bourigault et al., 2016; Tran and Zheleva, 2022). Propagation traces are data structures containing information about which entities were activated and when. Goyal et al. (2011a) uses these traces to create a model that does not make any parametric assumptions about the diffusion process, but instead learns the expected influence spread directly from data. These methods, however, do not consider node features. The only exception is Tran and Zheleva (2022), in which node features are leveraged to estimate the parameters in the Linear Threshold model, a commonly used diffusion process. In contrast to these approaches, our proposed method leverages node features, does not make parametric assumptions regarding the diffusion process, and does not use propagation traces. Moreover, OTAPI allows for out-of-sample IM, as the network for which we try to find the optimal treatment allocation may be different from the network that was used for training the model. This is an important advantage of OTAPI, as networks—especially social networks—typically change over time. Ma and Tresp (2021) proposes a utility maximization approach for instance-level treatment assignment in the presence of interference using a policy network that leverages a relational causal estimator. The budget constraint is satisfied by integrating it into the loss function of the policy network and allowing some tolerance for violating it. In comparison, OTAPI directly optimizes treatment allocation for the network as a whole by leveraging algorithms that can support hard constraints.
@@ 3 Problem setup
Notation.  Consider an undirected network $\mathcal{G}=(\mathcal{N},\mathcal{E})$ where $\mathcal{N}$ represents the set of nodes and $\mathcal{E}$ denotes the set of edges between them. We define the set of neighbors of node $i$ as $\mathcal{N}_{i}$, which comprises nodes $j\in\mathcal{N}$ for which an edge $\{i,j\}$ exists. In this context, $\{i,j\}=\{j,i\}$. Additionally, we denote the set of nodes not in the neighborhood of $i$ as $\mathcal{N}_{-i}=\mathcal{N}\setminus(\mathcal{N}_{i}\cup\{i\})$. Each node $i$ possesses a feature vector $\mathbf{X}_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$. We will use nodes and entities interchangeably, depending on the context.
Let $Y_{i}\in\mathcal{Y}$ represent the outcome of entity $i$. Without loss of generality, $\mathcal{Y}=\{0,1\}$, e.g., buying or not buying the product. We define $T_{i}\in\mathcal{T}=\{0,1\}$ as the treatment given to entity $i$. The treatments of $i$’s neighbors $\mathcal{N}_{i}$ is given by $\mathbf{T}_{\mathcal{N}_{i}}=\{T_{j}\}_{j\in\mathcal{N}_{i}}$. Following Forastiere et al. (2021), Jiang and Sun (2022), and Aronow and Samii (2017), an exposure mapping, denoted by $Z:\mathbf{T}_{\mathcal{N}_{i}}\to Z_{i}$, is defined to summarize $\mathbf{T}_{\mathcal{N}_{i}}$. This function describes how the outcome of an entity is dependent on $\mathbf{T}_{\mathcal{N}_{i}}$. The domain $\mathcal{Z}$ of the exposure mapping $Z_{i}$ depends on the exact definition of the function $Z$ (Forastiere et al., 2021). We follow the potential outcomes framework (Rubin, 1974), where $Y_{i}(T_{i}=t,Z_{i}=z)$ denotes the outcome value entity $i$ would take if it receives treatment $t$ with a neighborhood exposure $z$. A potential outcome is not necessarily observed. For readability, we will sometimes write $\mathbb{E}\left[Y_{i}(T_{i}=t,Z_{i}=z)\mid\mathbf{X}_{i}=\mathbf{x}_{i},\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}}\right]$ as $\mathbb{E}\left[Y_{i}(t,z)\mid\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}}\right]$.
@@
Figure 2: Graphical representation of the causal structure. Only two entities are included for simplicity. $Y_{i}$ might be influenced by its own features $\mathbf{X}_{i}$, the features of its neighbors $\mathbf{X}_{j}$, its own treatment $T_{i}$, and the treatment of its neighbors $T_{j}$. Moreover, $T_{i}$ might be influenced by both $\mathbf{X}_{i}$ and $\mathbf{X}_{j}$.
Assumptions.  We assume that an entity’s own treatment $T_{i}$, the treatments of its neighbors $\mathbf{T}_{\mathcal{N}_{i}}$, an entity’s features $\mathbf{X_{i}}$ and the features of its neighbors $\mathbf{X}_{\mathcal{N}_{i}}=\{\mathbf{X}_{j}\}_{j\in\mathcal{N}_{i}}$ might causally influence its outcome $Y_{i}$. The assumed causal structure is depicted in Figure 2 as a Directed Acyclic Graph (DAG). While it relaxes the “no interference" assumption, the DAG does not imply that we assume there always is interference. The presence of a causal arrow in this DAG implies that we do not assume anything about whether there truly is a causal effect (Pearl, 2009). Note that we do not consider contagion effects, meaning the outcomes of different entities do not causally affect each other (Ogburn and VanderWeele, 2014). To summarize $\mathbf{T}_{\mathcal{N}_{i}}$ we use the mapping $Z_{i}=\frac{\sum_{j\in\mathcal{N}_{i}}T_{j}}{|\mathcal{N}_{i}|}$ (Forastiere et al., 2021; Jiang and Sun, 2022) [^1] Note that more complex exposure mappings (e.g., multi-dimensional, based on node features, etc.) can also be defined/learned (Cai et al., 2023; Adhikari and Zheleva, 2023; Aronow and Samii, 2017).. We assume that the decision to treat is taken simultaneously for all entities. Consequently, both $Z_{i}$ and $T_{i}$ are part of the intervention on $i$.
Traditionally, each entity has two potential outcomes: $Y_{i}(T_{i}=0)$ and $Y_{i}(T_{i}=1)$. However, in the presence of interference, the traditional consistency assumption, which states that there are no multiple versions of treatment, is violated. Here, depending on the treatment of $i$’s neighbors, $i$’s potential outcomes might differ. To address this, the consistency assumption is relaxed by employing the exposure mapping $Z_{i}$ (Forastiere et al., 2021; Jiang and Sun, 2022):